OpenStudy (anonymous):

if g(x)=sqrt(x+1) and f(x)=(1/x^2-1) then how do you find the domain of (f o g)(x)

OpenStudy (anonymous):

find first (f o g)(x)

OpenStudy (anonymous):

(f o g)(x)=f(g(x))=1/(((sqrt(x+1))^2-1)=1/x

OpenStudy (anonymous):

so domain is R-{0}

OpenStudy (anonymous):